OpenStudy (diamondanon2):

Solve the following equation. Show your work. x^2 - 4 = 0.

jimthompson5910 (jim_thompson5910):

Hint: Difference of Squares Rule \[\Large a^2-b^2 = (a-b)(a+b)\]

OpenStudy (diamondanon2):

Sorry :( still lost

jimthompson5910 (jim_thompson5910):

So for example, if a = x and b = 3, then \[\Large a^2-b^2 = (a-b)(a+b)\] \[\Large x^2-3^2 = (x-3)(x+3)\] \[\Large x^2-9 = (x-3)(x+3)\]

jimthompson5910 (jim_thompson5910):

What would \(\LARGE x^2-4\) factor to?

OpenStudy (diamondanon2):

(x-2)(x+2)

jimthompson5910 (jim_thompson5910):

yes, so if x^2 - 4 = 0,then either x-2 = 0 or x+2 = 0

jimthompson5910 (jim_thompson5910):

so you need to solve x-2 = 0 and x+2 = 0 to get the two solutions

jimthompson5910 (jim_thompson5910):

sorry I skipped a step. I should have gone from x^2-4 = 0 to (x-2)(x+2) = 0 then break down to those two equations. But you get the idea hopefully

OpenStudy (diamondanon2):

so it would be x=2 and x=-2

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (diamondanon2):

thank you a bunch!

jimthompson5910 (jim_thompson5910):

Let's check each possible answer To check x = 2, we replace each x with 2 to get x^2-4 = 0 (2)^2-4 = 0 4-4 = 0 0 = 0 Similar work would be done to check x = -2

jimthompson5910 (jim_thompson5910):

no problem